OpenStudy (anonymous):

"Find the volume of the solid obtained by rotating the region bounded by the given curves about the given axis:" y = x^4, y = 1; about y = 2 I set up my integral as π∫(1 - x^4)^2 - (1)^2 dx from -1 to 1. I calculated this, several times, to get -(26π)/45. (I realize you take the absolute value of a negative area) The answer in the book is (208π)/45.

OpenStudy (anonymous):

why did u take (1-x^4)^2??

OpenStudy (anonymous):

and what is the limit in x direction??

OpenStudy (anonymous):

You are doing shell yet you are using washing equation

OpenStudy (anonymous):

I ended up figuring it out. I had to take (2 - x^4) as the radius since y = 2 was the middle of the whole big thing and x^4 was the bottom. The top minus the bottom then is the radius. Since y = x^4 and y = 1 bound the object, the limits in x direction are -1 and 1. You can use shells and washers interchangeably, just like you can solve for area in respect to both x and y to get the same value. This problem was in a section that dealt with washers as well. Shells won't be introduced till the next chapter.